# Multimodal 3D Reasoning Segmentation with Complex Scenes

### v2

@@ 4.4 Explainability
The LLM-generated textual answers $\hat{Y}_{ans}$ contain explanations of the implicit intention of the user questions. In addition, the answer includes the descriptions of 3D spatial relations of multiple objects which provide useful guidance for identifying these objects. Unlike most existing reasoning-based segmentation methods [38, 43, 13, 16] that just output statements like "It is <SEG>." without further explanation, our generated answer integrates detailed explanations about 3D spatial relations to enhance segmentation. The ground truth answers $Y_{ans}$ also incorporate these 3D spatial relations and are used to supervise the generated output textual answers, guiding the model to capture spatial information effectively. The textual answer loss is defined by:
@@ 4.5 Training Objectives
The proposed model is trained end-to-end with a textual answer loss $L_{ans}$ for textual answer generation, a mask loss $L_{mask}$ for point cloud segmentation mask prediction, and a classification loss $L_{cls}$ for point cloud classification. The overall loss function is formulated as follows:
@@ 5.1 Experimental Settings
3D Vision-Language Methods
@@
Implementation Details.
We conduct experiments on one NVIDIA V100 GPU and train the framework for 100 epochs with a batch size of 1. We use the Adam [22] optimizer with the initial learning rate of $1\times 10^{-4}$. We adopt LLaMA-7B [37] as our multimodal LLM backbone and LoRA [15] to perform efficient fine-tuning. All experiments are conducted on the proposed ReasonSeg3D dataset.
@@ 5.2 Benchmarking with Existing Methods
We benchmark MORE3D with state-of-the-art 3D segmentation and vision-language methods on the ReasonSeg3D validation set. As Table 2 shows, our method achieves superior reasoning segmentation performance across all evaluation metrics. Specifically, OpenScene [31], PLA [9], and RegionPLC [46] are designed to comprehend input words or phrases based on CLIP-based language models. They lack the ability to reason user intention from input questions and to correctly generate answers targeting at corresponding multiple objects. Consequently, the predicted 3D segmentation masks by these methods cannot be aligned correctly with ground-truth 3D segmentation masks, resulting in low cIoU and gIoU. Moreover, 3D-STMN [41] and RefMask3D [12] are tailored for single-object segmentation and therefore cannot handle multiple objects, while MDIN [40] exhibits weaker reasoning ability for interpreting implicit human intention. Additionally, 3D-LLM [14] and 3D-VisTA [53], designed for general 3D tasks, are re-implemented on our proposed 3D reasoning segmentation task for comparison. In contrast, MORE3D can accurately reason the implicit intention of user questions, predicting 3D segmentation masks that are more consistent with the ground truth.
@@
Figure 5 presents qualitative segmentation by the proposed MORE3D with two examples from the ReasonSeg3D validation set. Each example demonstrates the user question, input point cloud, ground-truth segmentation, and the 3D segmentation masks that are predicted by MORE3D. We can observe that MORE3D could accurately predict the 3D segmentation masks that are highly aligned with the ground-truth segmentation masks. Specifically, in the example at the top, despite many chairs closely surrounding the central long table, MORE3D can correctly distinguish and segment them. For the example at the bottom, despite the similar appearance and size of tables and chairs, MORE3D can distinguish and segment them precisely. The visualization indicates that MORE3D can comprehend the implicit intention behind the user questions and segment multiple objects accurately.
@@ 5.3 Ablation Study
We conduct extensive ablation studies on the ReasonSeg3D validation set to evaluate our designs. Specifically, we examine MORE3D from the aspect of the approach for object-specific point cloud embeddings extraction, the loss design, the point cloud decoding approach, as well as the way of prediction.
@@
We examine the effectiveness of the designed multi-seg index list $I_{seg}$ for the object-specific point cloud embeddings extraction in MORE3D. As Table 3 shows, using the multi-seg index list $I_{seg}$ can achieve better performance than adopting the random selection approach. This is largely because the LLM’s embeddings corresponding to multiple objects are extracted precisely with the help of $I_{seg}$. As a comparison, the random selection could extract embeddings unrelated to target objects like “table” or “chairs”, and it could also extract embeddings belonging to irrelevant non-object words, such as “around” and “suitable”. This explains why using the multi-seg index list $I_{seg}$ is more effective for extracting object-specific point cloud embeddings corresponding to multiple target objects.
@@
Table 3: Ablation study on the approach of object-specific point cloud embeddings extraction. The best results are in bold.
@@
Table 4: Ablation study of the loss functions on the ReasonSeg3D validation set. $L_{ans}$ and $L_{mask}$ refer to the textual answer loss and the point cloud segmentation mask loss, respectively. The best results are in bold.
@@
Decoding Operation.
Index Operation cIoU gIoU
1 Addition 28.26 31.01
2 Concatenation 29.42 31.17
3 Dot Product 30.19 32.01
Table 5: Ablation study on the approach of point cloud decoding operation in the 3D Decoder. The best results are in bold.
Index Manner cIoU gIoU
1 Unified 29.06 30.43
2 Separated 30.19 32.01
Table 6: Ablation study on different approaches of prediction in the 3D Decoder. The best results are in bold.
Prediction Approaches.
We examine the effectiveness of different prediction approaches including unified and separated approaches for the prediction of 3D segmentation masks and classification. Table 6 shows experimental results. For the unified approach, the 3D segmentation masks and classification results are predicted by a single instead of two separate network branches. This results in lower performance due to the increased learning burden when a single network branch must learn to perform segmentation and classification simultaneously. In contrast, the separate approach as illustrated in the bottom right of Figure 3 assigns each task to an independent network branch to learn. This allow each branch to focus on its respective task which lowers the learning burden and improves performance clearly.
@@ 6 Conclusion
This paper presents a novel multi-object 3D reasoning segmentation task, producing both 3D segmentation masks and textual explanations with rich 3D spatial relations among multiple objects within complex 3D scenes. To this end, we develop ReasonSeg3D, a large-scale benchmark featuring rich 3D spatial relations integrated into question-answer pairs, designed to evaluate multi-object 3D reasoning segmentation effectively. On top of this, we propose MORE3D, a technique for multi-object 3D reasoning segmentation with textual explanations, demonstrating strong reasoning abilities in response to user questions. Extensive experiments validate the effectiveness of MORE3D. Future work will focus on generalizing our work to diverse and challenging 3D environments, such as outdoor scenes, to expand the scope of applications.

### v3

@@ 4.4 Explainability
The LLM-generated textual answers $\hat{Y}_{ans}$ contain explanations of the implicit intention of the user’s questions. In addition, the answer includes the descriptions of 3D spatial relations of multiple objects which provide useful guidance for identifying these objects. Unlike most existing reasoning-based segmentation methods [39, 45, 13, 16] that just output statements like "It is <SEG>" without further explanation, our generated answer integrates detailed explanations about 3D spatial relations to enhance segmentation. The ground truth answers $Y_{ans}$ also incorporate these 3D spatial relations and are used to supervise the generated output textual answers, guiding the model to capture spatial information effectively. The textual answer loss is defined by:
@@ 4.5 Training Objectives
The proposed model is trained end-to-end with a textual answer loss $L_{ans}$ for textual answer generation, a mask loss $L_{mask}$ for point cloud segmentation mask prediction, and a classification loss $L_{cls}$ for point cloud classification. All three loss terms are assigned equal weights, and the overall loss function is formulated as follows:
@@ 5.1 Experimental Settings
Implementation Details.
We conduct experiments on one NVIDIA V100 GPU (32 GB memory) and train the network for 100 epochs with a batch size of 1, with a total training time of around 1 day. We use the Adam [23] optimizer with the initial learning rate of $1\times 10^{-4}$. We adopt LLaMA-7B [38] as our multimodal LLM backbone and LoRA [15] to perform efficient fine-tuning. All experiments are conducted on the proposed ReasonSeg3D dataset.
@@
         3D General Scene Understanding Methods
@@ 5.2 Benchmarking with Existing Methods
We benchmark MORE3D with state-of-the-art methods, including 3D segmentation [32, 9, 48, 43, 12, 42], general scene understanding [14, 58], and reasoning segmentation [16] approaches, on the ReasonSeg3D validation set. All compared methods are trained and evaluated on the same ReasonSeg3D to ensure a fair comparison. As Table 2 shows, our method achieves superior reasoning segmentation performance across all evaluation metrics. Specifically, OpenScene [32], PLA [9], and RegionPLC [48] are designed to comprehend input words or phrases based on CLIP-based language models. They lack the ability to reason user intention from input questions and to correctly generate answers targeting multiple corresponding objects. Consequently, the predicted 3D segmentation masks by these methods cannot be aligned correctly with ground-truth 3D segmentation masks, resulting in low cIoU and gIoU. Moreover, 3D-STMN [43] and RefMask3D [12] are tailored for single-object segmentation and therefore cannot handle multiple objects, while MDIN [42] exhibits weaker reasoning ability for interpreting implicit human intention. Additionally, 3D-LLM [14] and 3D-VisTA [58], designed for general 3D tasks, are re-implemented on our proposed 3D reasoning segmentation task for comparison. Furthermore, Reason3D [16], as a 3D reasoning segmentation approach, is limited to single-object settings and struggles with multi-object reasoning segmentation. In contrast, MORE3D can accurately reason the implicit intention of user questions, predicting 3D segmentation masks that are more consistent with the ground truth.
@@
Figure 5 presents qualitative segmentation by the proposed MORE3D with two examples from the ReasonSeg3D validation set. Each example demonstrates the user question, input point cloud, ground-truth segmentation, and the 3D segmentation masks that are predicted by MORE3D. We can observe that MORE3D could accurately predict the 3D segmentation masks that are highly aligned with the ground-truth segmentation masks. Specifically, in the top example, despite many chairs closely surrounding the central long table, MORE3D can correctly distinguish and segment them. In the bottom example, despite the similar appearance and size of tables and chairs, MORE3D can distinguish and segment them precisely. The visualization indicates that MORE3D can comprehend the implicit intention behind the user questions and segment multiple objects accurately.
@@ 5.3 Ablation Study
We conduct ablation studies on the ReasonSeg3D validation set to evaluate our designs. Specifically, we examine MORE3D from the aspect of the approach for object-specific embedding extraction, the way of prediction, the point cloud decoding approach, as well as the loss design.
@@
We examine the effectiveness of the designed multi-seg index list $I_{seg}$ for the object-specific embedding extraction in MORE3D. As Table 6 shows, using the multi-seg index list $I_{seg}$ can achieve better performance than adopting the random selection approach. This is largely because the LLM’s embeddings corresponding to multiple objects are extracted precisely with the help of $I_{seg}$. As a comparison, the random selection could extract embeddings unrelated to target objects like “table" or “chairs", and it could also extract embeddings belonging to irrelevant non-object words, such as “around" and “suitable". This explains why using the multi-seg index list $I_{seg}$ is more effective for extracting object-specific point cloud embeddings corresponding to multiple target objects.
Table 3: Ablation study on object-specific embedding extraction approach.
@@
Table 4: Ablation study on different approaches of prediction in the 3D Decoder.
Index Manner cIoU gIoU
1 Unified 29.06 30.43
2 Separated 30.19 32.01
Table 5: Ablation study on the approach of point cloud decoding operation in the 3D Decoder.
Index Operation cIoU gIoU
1 Addition 28.26 31.01
2 Concatenation 29.42 31.17
3 Dot Product 30.19 32.01
Table 6: Ablation study of the loss functions on ReasonSeg3D validation set.
@@
Prediction Approaches.
We examine the effectiveness of different prediction approaches including unified and separated approaches for the prediction of 3D segmentation masks and classification. Table 6 shows experimental results. For the unified approach, the 3D segmentation masks and classification results are predicted by a single network instead of two separate network branches. This results in lower performance due to the increased learning burden when a single network branch must learn to perform segmentation and classification simultaneously. In contrast, the separate approach as illustrated in the bottom right of Figure 3 assigns each task to an independent network branch to learn. This allows each branch to focus on its respective task which lowers the learning burden and improves performance clearly.
Decoding Operation.
@@ 6 Conclusion
This paper presents a novel multi-object 3D reasoning segmentation task, producing both 3D segmentation masks and textual explanations with rich 3D spatial relations among multiple objects within complex 3D scenes. To this end, we develop ReasonSeg3D, a large-scale benchmark featuring rich 3D spatial relations integrated into question-answer pairs, designed to evaluate multi-object 3D reasoning segmentation effectively. On top of this, we propose MORE3D, a technique for multi-object 3D reasoning segmentation with textual explanations, demonstrating strong reasoning abilities in response to user questions. Extensive experiments validate the effectiveness of MORE3D. Future work will focus on exploring our work in more challenging 3D environments to improve applicability in real-world scenarios.